3) In the space before each of the following transactions, identify if the transaction is an:

           Asset Source Transaction

          Asset Use Transaction

            Asset Exchange Transaction

__ ___   a.) Acquired $80,000 to start a business by issuing common stock.

__ ___   b.) Earned $18,000 of Revenue on account.

__ ___   c.) Paid $12,000 for one year’s rent in advance.

__ ___   d.) Purchased $8,000 of supplies on account.

__ ___   e.) Recognized $1,000 of rent expense used during the accounting period. (Previously

                        paid for in Transaction   c.)    above).

__ ___   f.) Paid $5,000 to suppliers in partial settlement of Accounts Payable.

4) Record each of the 6 transactions in Question #3 on the Horizontal Financial Statement on the next page. Be sure to calculate the totals in the last row of the chart to show whether the equations are correct.

Horizontal Financial Statement

Balance Sheet

              Income Statement

Cashflow Statement

                             Assets

=

       Liabilities

+

       Stockholders Equity

  

OA IA FA

Cash

A/R

Pre-Paid Rent

Supplies

=

A/P

N/P

+

Comm. Stk

R.E.

Rev

-

Exp

=

Net Inc

Event

=

+

-

=

a

=

+

-

=
Answer #1

Answer

--Requirement, as asked

Balance Sheet               Income Statement Cashflow Statement
                             Assets =        Liabilities +        Stockholders Equity    OA IA FA
Cash A/R Pre-Paid Rent Supplies = A/P N/P + Comm. Stk R.E. Rev - Exp = Net Inc
Event = + - =
a $80,000 = + $80,000 - = $80,000 FA

b

$18,000 $18,000 $18,000 $18,000
c ($12,000) $12,000 ($12,000) OA
d $8,000 $8,000
e ($1,000) ($1,000) $1,000 ($1,000)
f ($5,000) ($5,000) ($5,000) OA
TOTAL $63,000 $18,000 $11,000 $8,000 $0 $3,000 $0 $0 $0 $80,000 $17,000 $0 $18,000 $0 $1,000 $0 $17,000 $63,000 NC